Problem 10A-8 Applying Overhead; Overhead Variances (LO10-3, LO10-4] Lane Company manufactures a single product that requiresefficiency variances and the fixed overhead budget and volume variances. Complete this question by entering your answers in tReq 1 Reg 2 Req ЗА Req 3B Reg 4 Prepare a standard cost card for the companys product. (Round your answers to 2 decimal placReq1 Req 1 Reqz Req 3A Req 3A Req B Reg 4 Compute the standard direct labor-hours allowed for the years production. StandardReq 1 Req 2 Req ЗА Req 3B Req 4 Complete the following Manufacturing Overhead T-account for the year. Manufacturing OverheadReg 1 Reg 2 Req ЗА Req 3B Req 4 Determine the reason for any underapplied or overapplied overhead for the year by computing t
Answer #1

Solution 1:

Predetermined overhead rate for the year = Budgeted overhead / Budgeted labor hours

= [(105000*$3) + $735,000) / 105000 = $10 per direct labor hour

Variable overhead rate per labor hour = $3

Fixed overhead rate per labor hour = $7

Solution 2:

Standard cost Card - Lane Company
Particulars Qty Rate Per unit
Direct material 4 Pound at $5.50 Per Pound $22.00
Direct labor 1.5 DLHs at $12.50 Per DLH $18.75
Variable overhead 1.5 DLHs at $3.00 Per DLH $4.50
Fixed overhead 1.5 DLHs at $7.00 Per DLH $10.50
Standard cost per unit $55.75

Solution 3a:

standard direct labor-hours allowed for the year’s production = 84000*1.50 = 126000 hours

Solution 3b:

Manufacturing overhead - Lane Company
Particulars Debit Particulars Credit
Variable overhead incurred $259,350.00 Applied overhead (84000*1.5*$10) $1,260,000.00
Fixed overhead incurred $750,750.00
Overapplied overhead $249,900.00
Total $1,260,000.00 Total $1,260,000.00

Solution 4:

Actual rate of variable overhead = $259,350 / 136500 = $1.90 per labor hour

Variable overhead rate variance = (SR - AR) * AH = ($3 - $1.90) * 136500 = $150,150 F

Variable overhead efficiency variance = (SH - AH) *SR = (126000 - 136500) * $3 = $31,500 U

Fixed overhead budget variance = Budgeted fixed overhead - Actual fixed overhead = $735,000 - $750,750 = $15,750 U

Fixed overhead volume variance = Fixed overhead applied - Budgeted fixed overhead = (126000*$7) - $735,000

= $147,000 F